# How Do AI Agents Do Human Work? Comparing AI and Human Workflows Across Diverse Occupations

### v1

@@ 1 Introduction
AI agents are increasingly developed to perform tasks traditionally carried out by human workers (Brynjolfsson et al., 2018; Eloundou et al., 2023; Bick et al., 2024), as reflected in the growing competence of computer-use agents in work-related tasks such as software engineering (Jimenez et al., 2024; Wang et al., 2025a) and writing (Clark et al., 2018). Nonetheless, they still face challenges in many scenarios such as basic administrative or open-ended design tasks (Xu et al., 2024), sometimes creating a gap between expectations and reality in agent capabilities to perform real-world work (Zhang et al., 2021; Ryoo et al., 2025).
@@
To further improve agents’ utility at such tasks, we argue that it is necessary to look beyond their end-task outcome evaluation as measured in existing studies (Patwardhan et al., 2025), and investigate how agents currently perform human work — understanding their underlying workflows to gain deeper insights into their work process, especially how it aligns or diverges from human workers, to reveal the distinct strengths and limitations between them. Therefore, such an analysis should not benchmark agents in isolation (Xu et al., 2024; Patwardhan et al., 2025), but rather be grounded in comparative studies of human and agent workflows. More broadly speaking, to foster a collaborative future where humans and AI work effectively together, it is crucial to understand how both operate in practice, and to anticipate the organizational and behavioral changes required as agents assume greater roles (Masters et al., 2025).
@@
To this end, we first propose to perform a direct comparison of how human and agent workers do the same tasks. Beyond specific domains (Jimenez et al., 2024; Choi and Schwarcz, 2024), we aim to provide a holistic view of how agents might reshape the broader human workforce, by systematically studying human and agent workflows across occupations via essential shared skills — data analysis, engineering, computation, writing, and design — instantiated with 16 realistic, long-horizon tasks. We estimate these tasks to be representative of 287 computer-using U.S. occupations and 71.9% of their daily work activities (§2). Comparing human and agent activities is not trivial: we lack a unified representation and a scalable channel for analysis. We first collect human activities in a similar way to agent activities with actions and screen states. However, activities in their raw format, low-level mouse and keyboard actions, are largely context-independent (i.e., a single key press may not map to a meaningful user-intended task) (Shaw et al., 2023) thus unintuitive to interpret for both human and agent readers (Wang et al., 2024b). Inspired by humans who naturally reason about work in terms of higher-level workflows: structured sequences of actions performed together to achieve a goal (Deka et al., 2016; Liu et al., 2018; Wang et al., 2025c). We develop a workflow induction procedure that transforms heterogeneous computer-use activities into hierarchical, interpretable workflows. These workflows provide a shared representation that enables effective comparative analysis in our study (§3).
@@ 2.1 Representative Work-Related Tasks
Building the Skill Taxonomy  We manually identify shared skills from the list of task requirements across computer-related occupations. More concretely, we source all occupation records from the U.S. Department of Labor’s O*NET database (National Center for O*NET Development, 2024), which contains 923 occupations with 18,796 task requirements in the U.S. labor market. As our focus is on computer-using tasks that may be assisted or automated by digital agents, we filter tasks following Shao et al. (2025) and yield 2131 tasks from 287 occupations. We annotate the high-level skills required for tasks, and focus on skills that are feasible for automation. In the end, we identified five major skill categories (data analysis, engineering, computation, writing, and design) that, in aggregate, affect these 287 computer-using U.S. occupations and 71.9% of the daily tasks involved, according to the O*NET records. This high coverage ensures that our study can affect a large portion of human work. More details on skill annotation are in §A.1.
@@
To benchmark complex, realistic work-related tasks, we follow TheAgentCompany (TAC) (Xu et al., 2024) setup, where each instance contains a task instruction, necessary environmental contexts (e.g., input files, software at pre-work states), and an executable program evaluator for rigorous correctness checking. We adopt TAC’s multi-checkpoint evaluation protocol to capture progress at intermediate stages, enabling fine-grained assessment beyond the final outcomes. For reproducibility and fair comparison, all experiments run in TAC’s sandboxed environments, which host engineering tools (bash, python) and work-related websites for file sharing, communication, and project management. We adopt TAC tasks for skills such as data analysis and engineering. But because TAC is based on a software engineering company, we need to create additional tasks, particularly writing and design, to better represent occupations across all skill domains. Refer to §A.2 for the full list of tasks we studied.
@@ 2.2 Collecting Human Worker Activities
Processing Raw Computer-Use Activities  Computer-use activities collected from human real-world usage can be noisy and redundant, due to the limitations in the recording tool implementation. We hence perform post-processing to merge consecutive keypress and scroll actions, as well as merging two click() within 0.1 seconds into a single double_click(). This effectively simplifies the trajectory data, reducing the action count by 83.2% (from 5831.1 to 981.1 actions) in an average human trajectory, meanwhile aligning better with the granularity of agent computer activities (Wang et al., 2025b).
@@ 2.3 Collecting AI Agent Activities
In summary, we collected 112 trajectories for 16 tasks spanning across the 5 major skill categories. We collected 64 agent and 48 human trajectories, which constitute 33.8 and 981.1 steps on average, [^2] Human actions are often taken on a finer granularity (e.g., key_press({a short phrase})) compared to agent actions (e.g., key_press({an entire file})) and contain more noisy shifts (e.g., between type and click). substantially higher than previous computer-use agent benchmarks (e.g., the best agents take 5.9 steps to solve an average WebArena task (Wang et al., 2025c)), signaling the greater complexity and realism of the work-related tasks we study.
@@ 4.1 The Human Way, The Agent Way
To concretely test this hypothesis, we measure the alignment of agent workflows to human workflows that use programmatic tools and non-programmatic tools, respectively. Since the high-level steps mostly match (e.g., human step 4 and agent step 3 in Figure 3), we analyze steps one level deeper in the workflow hierarchy (e.g., finer steps in Figure 3, a). As shown in Figure 4 (c), agents exhibit 27.8% stronger alignment with program-using human steps than those that do not. This highlights a defining characteristic of agent workers — a programmatic mode of operation. This is not only true for coding-oriented OpenHands agents, but also for general-purpose computer-use agents such as ChatGPT and Manus. While this programmatic tendency is understandable in domains represented in LM code training data (e.g., engineering), this programmatic behavior notably persists even in visually intensive tasks such as slide creation and graphical design. Not only is such ‘blind’ visual creation possible, but LM-supported agents may in fact be more proficient at editing in the symbolic space (i.e., write programs) than in the visual space (i.e., adjust pixels) (Qiu et al., 2024).
This stronger alignment among workflows using similar tools also resonates with the notion of tool affordance from a design perspective (Norman, 2013)We shape our tools, and thereafter our tools shape us (McLuhan, 1994). When two workers, whether human or agent, employ the same tool to complete the same workflow step, their workflows tend to converge more closely at finer levels of granularity. Moreover, the design of human-agent interfaces must account for the granularity of interaction (Liang et al., 2025): when collaboration occurs at high-level workflow steps, it may suffice for a tool to be accessible to either the agent or the human; at finer granularities, however, the tool must support both programmatic and UI-based actions for seamless coordination (Song et al., 2024).
@@ 4.2 Human Workflows are Substantially Altered by AI Automation
How does Using AI Tools Affect Human Workflows?  Among human workers who use AI tools, the majority, 75.0% in our study, do so for augmentation purposes (Handa et al., 2025). Compared to those who complete tasks independently (without AI assistance; e.g., explore Figma templates to find logo design insights), humans using AI for augmentation (delegate a specific step; e.g., ask ChatGPT for design insights) show minimal changes in their overall workflows. In such cases, AI functions as an alternative tool for an existing step, similar to other tool choices in Figure 11. In contrast, workers employing AI tools for automation (rely on AI-driven workflows for entire tasks) experience substantial workflow changes: their activities shift from hands-on ‘building’ to ‘reviewing’ or ‘debugging’ AI-drafted solutions. One example is shown in Figure 5, where extra review steps emerge throughout a data processing workflow.
@@
We also find that, relative to independent workers who do not use AI, AI augmentation accelerates human work by 24.3%, whereas AI automation slows humans down by 17.7%. This slowdown can be partly attributed to two factors: (i) additional time spent verifying, debugging, and correcting AI solutions, often exceeding the time required to perform the task manually (Becker et al., 2025); and (ii) lower task expertise among human workers who are more likely to rely on AI for full-process automation rather than selective step-level assistance. Also, extra time (i) may be more likely to be incurred by less experienced workers (ii) when completing the task.
@@ 5.1 Agents Fail in Concerning Ways
$\bullet$ Formatting  One such example involves formatting (Figure 6, c), such as the use of special symbols and color schemes in data sheets, or refined font and layout choices in written documents. Visually polished formatting for varied work-related outcomes, such as docs or slides, conveys professionalism and fosters trust in a worker’s expertise (Tractinsky et al., 2000; Fogg et al., 2001), thereby enhancing the credibility of the delivered work (Li et al., 2018). While such formatting is straightforward and intuitive through UI interfaces, it is understandably neglected by agents that generate outputs programmatically without visual feedback or rendering.
@@ 6.1 Workflow-Inspired Agent Designs
Human Workflows As Demonstrations  One direct way to improve agents is to draw upon human expert workflows (Cypher, 1991). This is especially beneficial for tasks that are not naturally programmable, such as parsing data from bill images into an Excel file. As discussed above, these tasks are challenging for agents because they lack clear programming solutions. In the Figure 6 (a) example, the agent uses an unstructured approach and fails to extract bill data accurately. When human workflows induced by our tool (§3) are augmented as contexts to the agent, the agent begins to emulate the human procedure: viewing and extracting data from bills one-by-one, and finally aggregating them into an Excel file, correctly solving the task this time.
@@
Workflow Elaboration During Task-Solving  For tasks that agents execute the full workflow but fail to deliver correct outcomes, it would be valuable for humans to pinpoint specific steps where errors emerge, and intervene before agents fabricate deliverables (Horvitz, 1999). However, we find this challenging in practice, as the agent rarely verbalizes their deviations or errors. Consequently, the induced workflow steps often merely restate the user’s original instruction (e.g., “create a Python script to extract data from 5 receipts just viewed”), despite that the agent acts differently (e.g., makes up 5 data entries and exports them to Excel). One potential way is to train agents to explicitly disclose such dishonest behaviors. Alternatively, the workflow induction process could be adapted to a supervisory angle to detect misalignment between the intended instruction and the actual actions taken by the agent.
Once such fault-informative workflows can be derived, especially on the fly as agents solve tasks, it would allow humans to provide more timely and accurate guidance (Figure 8). This can (i) help novices learn essential skills by observing agent workflows, and (ii) allow experts to exercise scalable oversight through more efficient verification and decision-making processes (Grunde-McLaughlin et al., 2025).
@@ 6.2 Towards More Capable Agents
While programming tools dominate agent workflows across domains (§4.1), the majority of current agent development efforts remain focused on software engineering applications (Jimenez et al., 2024; Pan et al., 2024; Jain et al., 2025). Although only 12.2% of occupations are engineering-focused (§2.1), programming skills are required in 82.5% of occupations — indicating that 60.3% of programming use cases serve non-engineering purposes. We therefore call for greater attention to building programmatic agents for non-engineering tasks (Wang et al., 2024b; Ge et al., 2025; Soni et al., 2025), for both training agents in versatile programming scenarios, and developing programmatic tools to support varied functionalities. After all, programs can be the hammer, not just the nail.
Stronger Vision Capabilities  Our analysis reveals that current agents have limited visual abilities. While most progress in visual understanding focuses on natural scenes (Deng et al., 2009), emerging efforts are addressing digital visual contexts (Gou et al., 2024; Xie et al., 2025). Strengthening agent’s foundational visual and UI-interaction skills through larger-scale training is essential, as programmatic workarounds are not always feasible for agents.
For visually intensive tasks such as logo or web design, enhancing agents’ visual perception alone may be insufficient. Programming-oriented agents often find it easier to edit visuals symbolically (i.e., with code) than to manipulate them directly on a canvas (Ma and Agrawala, 2025). As human-agent collaboration in visual work continues, rethinking interface designs and communication granularity will be key to bridging programmatic and UI-based worklows.
More Accurate and Holistic Work Quality Evaluations  More accurate and holistic evaluation frameworks are needed to assess agents at work on long-horizon and realistic tasks. Task success is not always singularly defined, as work can be correct in multiple ways, yet current program verifiers may not adequately capture the full spectrum of valid procedures and outcomes (Patwardhan et al., 2025).
Beyond the dimensions we studied (workflows, tool affordance, efficiency, and quality), many factors can shape how a worker’s ability and trustworthiness are perceived, including creativity and practicality (Mumford and Gustafson, 1988), communication (Shao et al., 2025) and adaptability (Pulakos et al., 2000) during the process, among other meta-aspects of work such as robustness (Barrick and Mount, 1991) and ethical behaviors (Treviño et al., 2006). Identifying and systematically studying these overlooked dimensions will be key to developing agents truly prepared for integration into the human workforce.

### v2

@@ 1 Introduction
AI agents are increasingly developed to perform tasks traditionally carried out by human workers (brynjolfsson2018what; eloundou2023gpts; bick2024the), as reflected in the growing competence of computer-use agents in work-related tasks such as software engineering (jimenez2024swebench; wang2025openhands) and writing (clark2018creative). Nonetheless, they still face challenges in many scenarios such as basic administrative or open-ended design tasks (xu2024theagentcompany), sometimes creating a gap between expectations and reality in agent capabilities to perform real-world work (zhang2021ideal; ryoo2025high).
@@
To further improve agents’ utility at such tasks, we argue that it is necessary to look beyond their end-task outcome evaluation as measured in existing studies (patwardhan2025gdpval), and investigate how agents currently perform human work — understanding their underlying workflows to gain deeper insights into their work process, especially how it aligns or diverges from human workers, to reveal the distinct strengths and limitations between them. Therefore, such an analysis should not benchmark agents in isolation (xu2024theagentcompany; patwardhan2025gdpval), but rather be grounded in comparative studies of human and agent workflows. More broadly speaking, to foster a collaborative future where humans and AI work effectively together, it is crucial to understand how both operate in practice, and to anticipate the organizational and behavioral changes required as agents assume greater roles (masters2025orchestrating).
@@
To this end, we first propose to perform a direct comparison of how human and agent workers do the same tasks. Beyond specific domains (jimenez2024swebench; choi2024ai), we aim to provide a holistic view of how agents might reshape the broader human workforce, by systematically studying human and agent workflows across occupations via essential shared skills — data analysis, engineering, computation, writing, and design — instantiated with 16 realistic, long-horizon tasks. We estimate these tasks to be representative of 287 computer-using U.S. occupations and 71.9% of their daily work activities (§2). Comparing human and agent activities is not trivial: we lack a unified representation and a scalable channel for analysis. We first collect human activities in a similar way to agent activities with actions and screen states. However, activities in their raw format, low-level mouse and keyboard actions, are largely context-independent (i.e., a single key press may not map to a meaningful user-intended task) (shaw2023from) thus unintuitive to interpret for both human and agent readers (wang2024trove). Inspired by humans who naturally reason about work in terms of higher-level workflows: structured sequences of actions performed together to achieve a goal (deka2016erica; zheran2018reinforcement; wang2025agent). We develop a workflow induction procedure that transforms heterogeneous computer-use activities into hierarchical, interpretable workflows. These workflows provide a shared representation that enables effective comparative analysis in our study (§3).
@@ 2.1 Representative Work-Related Tasks
Building the Skill Taxonomy  We manually identify shared skills from the list of task requirements across computer-related occupations. More concretely, we source all occupation records from the U.S. Department of Labor’s O*NET database (onet2024), which contains 923 occupations with 18,796 task requirements in the U.S. labor market. As our focus is on computer-using tasks that may be assisted or automated by digital agents, we filter tasks following shao2025future and yield 2131 tasks from 287 occupations. We annotate the high-level skills required for tasks, and focus on skills that are feasible for automation. In the end, we identified five major skill categories (data analysis, engineering, computation, writing, and design) that, in aggregate, affect these 287 computer-using U.S. occupations and 71.9% of the daily tasks involved, according to the O*NET records. This high coverage ensures that our study can affect a large portion of human work. More details on skill annotation are in §A.1.
@@
To benchmark complex, realistic work-related tasks, we follow TheAgentCompany (TAC) (xu2024theagentcompany) setup, where each instance contains a task instruction, necessary environmental contexts (e.g., input files, software at pre-work states), and an executable program evaluator for rigorous correctness checking. We adopt TAC’s multi-checkpoint evaluation protocol to capture progress at intermediate stages, enabling fine-grained assessment beyond the final outcomes. For reproducibility and fair comparison, all experiments run in TAC’s sandboxed environments, which host engineering tools (bash, python) and work-related websites for file sharing, communication, and project management. We adopt TAC tasks for skills such as data analysis and engineering. But because TAC is based on a software engineering company, we need to create additional tasks, particularly writing and design, to better represent occupations across all skill domains. Refer to §A.2 for the full list of tasks we studied.
@@ 2.2 Collecting Human Worker Activities
Processing Raw Computer-Use Activities  Computer-use activities collected from human real-world usage can be noisy and redundant, due to the limitations in the recording tool implementation. We hence perform post-processing to merge consecutive keypress and scroll actions, as well as merging two click() within 0.1 seconds into a single double_click(). This effectively simplifies the trajectory data, reducing the action count by 83.2% (from 5831.1 to 981.1 actions) in an average human trajectory, meanwhile aligning better with the granularity of agent computer activities (wang2025opencua).
@@ 2.3 Collecting AI Agent Activities
In summary, we collected 112 trajectories for 16 tasks spanning across the 5 major skill categories. We collected 64 agent and 48 human trajectories, which constitute 33.8 and 981.1 steps on average, [^2] Human actions are often taken on a finer granularity (e.g., key_press({a short phrase})) compared to agent actions (e.g., key_press({an entire file})) and contain more noisy shifts (e.g., between type and click). substantially higher than previous computer-use agent benchmarks (e.g., the best agents take 5.9 steps to solve an average WebArena task (wang2025agent)), signaling the greater complexity and realism of the work-related tasks we study.
@@ 4.1 The Human Way, The Agent Way
To concretely test this hypothesis, we measure the alignment of agent workflows to human workflows that use programmatic tools and non-programmatic tools, respectively. Since the high-level steps mostly match (e.g., human step 4 and agent step 3 in Figure 3), we analyze steps one level deeper in the workflow hierarchy (e.g., finer steps in Figure 3, a). As shown in Figure 4 (c), agents exhibit 27.8% stronger alignment with program-using human steps than those that do not. This highlights a defining characteristic of agent workers — a programmatic mode of operation. This is not only true for coding-oriented OpenHands agents, but also for general-purpose computer-use agents such as ChatGPT and Manus. While this programmatic tendency is understandable in domains represented in LM code training data (e.g., engineering), this programmatic behavior notably persists even in visually intensive tasks such as slide creation and graphical design. Not only is such ‘blind’ visual creation possible, but LM-supported agents may in fact be more proficient at editing in the symbolic space (i.e., write programs) than in the visual space (i.e., adjust pixels) (qiu2024can).
This stronger alignment among workflows using similar tools also resonates with the notion of tool affordance from a design perspective (norman2013design)We shape our tools, and thereafter our tools shape us (mcluhan1994understanding). When two workers, whether human or agent, employ the same tool to complete the same workflow step, their workflows tend to converge more closely at finer levels of granularity. Moreover, the design of human-agent interfaces must account for the granularity of interaction (liang2025agentbuilder): when collaboration occurs at high-level workflow steps, it may suffice for a tool to be accessible to either the agent or the human; at finer granularities, however, the tool must support both programmatic and UI-based actions for seamless coordination (song2024beyond).
@@ 4.2 Human Workflows are Substantially Altered by AI Automation
How does Using AI Tools Affect Human Workflows?  Among human workers who use AI tools, the majority, 75.0% in our study, do so for augmentation purposes (handa2025economic). Compared to those who complete tasks independently (without AI assistance; e.g., explore Figma templates to find logo design insights), humans using AI for augmentation (delegate a specific step; e.g., ask ChatGPT for design insights) show minimal changes in their overall workflows. In such cases, AI functions as an alternative tool for an existing step, similar to other tool choices in Figure 11. In contrast, workers employing AI tools for automation (rely on AI-driven workflows for entire tasks) experience substantial workflow changes: their activities shift from hands-on ‘building’ to ‘reviewing’ or ‘debugging’ AI-drafted solutions. One example is shown in Figure 5, where extra review steps emerge throughout a data processing workflow.
@@
We also find that, relative to independent workers who do not use AI, AI augmentation accelerates human work by 24.3%, whereas AI automation slows humans down by 17.7%. This slowdown can be partly attributed to two factors: (i) additional time spent verifying, debugging, and correcting AI solutions, often exceeding the time required to perform the task manually (becker2025measuring); and (ii) lower task expertise among human workers who are more likely to rely on AI for full-process automation rather than selective step-level assistance. Also, extra time (i) may be more likely to be incurred by less experienced workers (ii) when completing the task.
@@ 5.1 Agents Fail in Concerning Ways
$\bullet$ Formatting  One such example involves formatting (Figure 6, c), such as the use of special symbols and color schemes in data sheets, or refined font and layout choices in written documents. Visually polished formatting for varied work-related outcomes, such as docs or slides, conveys professionalism and fosters trust in a worker’s expertise (tractinsky2000beautiful; fogg2001makes), thereby enhancing the credibility of the delivered work (li2018communicating). While such formatting is straightforward and intuitive through UI interfaces, it is understandably neglected by agents that generate outputs programmatically without visual feedback or rendering.
@@ 6.1 Workflow-Inspired Agent Designs
Human Workflows As Demonstrations  One direct way to improve agents is to draw upon human expert workflows (cypher1991eager). This is especially beneficial for tasks that are not naturally programmable, such as parsing data from bill images into an Excel file. As discussed above, these tasks are challenging for agents because they lack clear programming solutions. In the Figure 6 (a) example, the agent uses an unstructured approach and fails to extract bill data accurately. When human workflows induced by our tool (§3) are augmented as contexts to the agent, the agent begins to emulate the human procedure: viewing and extracting data from bills one-by-one, and finally aggregating them into an Excel file, correctly solving the task this time.
@@
Workflow Elaboration During Task-Solving  For tasks that agents execute the full workflow but fail to deliver correct outcomes, it would be valuable for humans to pinpoint specific steps where errors emerge, and intervene before agents fabricate deliverables (horvitz1999principles). However, we find this challenging in practice, as the agent rarely verbalizes their deviations or errors. Consequently, the induced workflow steps often merely restate the user’s original instruction (e.g., “create a Python script to extract data from 5 receipts just viewed”), despite that the agent acts differently (e.g., makes up 5 data entries and exports them to Excel). One potential way is to train agents to explicitly disclose such dishonest behaviors. Alternatively, the workflow induction process could be adapted to a supervisory angle to detect misalignment between the intended instruction and the actual actions taken by the agent.
Once such fault-informative workflows can be derived, especially on the fly as agents solve tasks, it would allow humans to provide more timely and accurate guidance (Figure 8). This can (i) help novices learn essential skills by observing agent workflows, and (ii) allow experts to exercise scalable oversight through more efficient verification and decision-making processes (grunde2025designing).
@@ 6.2 Towards More Capable Agents
While programming tools dominate agent workflows across domains (§4.1), the majority of current agent development efforts remain focused on software engineering applications (jimenez2024swebench; pan2024training; jain2025regym). Although only 12.2% of occupations are engineering-focused (§2.1), programming skills are required in 82.5% of occupations — indicating that 60.3% of programming use cases serve non-engineering purposes. We therefore call for greater attention to building programmatic agents for non-engineering tasks (wang2024trove; ge2025autopresent; soni2025coding), for both training agents in versatile programming scenarios, and developing programmatic tools to support varied functionalities. After all, programs can be the hammer, not just the nail.
Stronger Vision Capabilities  Our analysis reveals that current agents have limited visual abilities. While most progress in visual understanding focuses on natural scenes (deng2009imagenet), emerging efforts are addressing digital visual contexts (gou2024navigating; xie2025scaling). Strengthening agent’s foundational visual and UI-interaction skills through larger-scale training is essential, as programmatic workarounds are not always feasible for agents.
For visually intensive tasks such as logo or web design, enhancing agents’ visual perception alone may be insufficient. Programming-oriented agents often find it easier to edit visuals symbolically (i.e., with code) than to manipulate them directly on a canvas (ma2025mover). As human-agent collaboration in visual work continues, rethinking interface designs and communication granularity will be key to bridging programmatic and UI-based worklows.
More Accurate and Holistic Work Quality Evaluations  More accurate and holistic evaluation frameworks are needed to assess agents at work on long-horizon and realistic tasks. Task success is not always singularly defined, as work can be correct in multiple ways, yet current program verifiers may not adequately capture the full spectrum of valid procedures and outcomes (patwardhan2025gdpval).
Beyond the dimensions we studied (workflows, tool affordance, efficiency, and quality), many factors can shape how a worker’s ability and trustworthiness are perceived, including creativity and practicality (Mumford1988CreativitySI), communication (shao2025future) and adaptability (pulakos2000adaptability) during the process, among other meta-aspects of work such as robustness (Barrick1991THEBF) and ethical behaviors (Trevio2006BehavioralEI). Identifying and systematically studying these overlooked dimensions will be key to developing agents truly prepared for integration into the human workforce.